# ISAC Network Planning: Sensing Coverage Analysis and 3-D BS Deployment Optimization

### v1

@@ II-B Cooperative Sensing Model
We aim to explore the optimal BS deployment method by examining time-of-flight based ranging measurements. The location of a certain target point is denoted as ${\bf{t}}$, where ${\bf{t}}\in\mathcal{A}$, and $\mathcal{A}$ represents the entire area of interest. The base-band equivalent of the signal reflected from target point ${\bf{t}}$ at receiver $m$ is represented as
@@
where $\beta\geq 2$ is the pathloss exponent between the serving BS and the target. [^3] For sensing performance analysis, we consider an open area of interest for target sensing where no blockage exists. This is because the radar sensing normally focuses on the light-of-sight (LoS) targets parameters and Non-LoS (NLoS) reflections may be neglected due to additional high pathloss of double-bouncing nature of radar. Furthermore, $\sigma$ denotes the radar cross section (RCS), $\tau_{n,m}$ is the propagation delay of the bistatic link spanning from BS $n$ to the target and then to BS $m$, while $\tilde{\tau}_{n,m}$ denotes the propagation delay of the direct link impinging from BS $n$ to BS $m$. In (3), ${\bf{H}}_{n,m}$ denotes the channel from BS $n$ to BS $m$. Finally, the term ${\bf{n}}(\tau)$ is the additive complex Gaussian noise having zero mean and covariance matrix $\sigma_{s}^{2}{\bf{I}}_{M_{\rm{r}}}$. In (3), we have $\bm{a}_{M_{\mathrm{t}}}\left(\Omega_{n}\right)=\left[1,\cdots,e^{{-j\pi\left(M_{\mathrm{t}}-1\right)\Omega_{n}}}\right]^{T}$.
@@
${\rm{var}}\{\hat{{\bf{t}}}\}={\rm{E}}\{|\hat{{\bf{t}}}-{{\bf{t}}}|^{2}\}\geq\mathrm{CRLB}\left({{\bf{t}}}\right),\vspace{-1.5mm}$ (4)
where $\hat{{\bf{t}}}=\left[\hat{x}^{t},\hat{y}^{t},\hat{z}^{t}\right]^{T}$ represents the estimated location of the target. By applying matched filtering and reciprocal filtering, one can estimate the ToF of the signal between targets and BSs, and hence determine the propagation distance. Specifically, from transmitter $n$ to the target and then to receiver $m$, the term $\hat{d}_{nm}$ represents the distance measured between the $m$th transmitter and the $n$th receiver, which is given by
$\hat{d}_{nm}=d_{nm}+n^{t}_{ij},\vspace{-1.5mm}$ (5)
where $d_{nm}$ denotes the true distance, the measurement noise is $n^{t}_{i,j}\sim\mathcal{N}\left(0,\eta_{i,j}^{2}\right)$ and $\eta_{ij}^{2}=\frac{3c^{2}\sigma_{s}^{2}}{8\pi^{2}G_{t}M_{r}B^{2}\gamma_{ij}}$. Here, $c$ denotes the speed of light, $B^{2}$ represents the effective squared bandwidth, $G_{t}$ is the transmit beamforming gain in the direction of the target, and $\gamma_{ij}$ represents the bistatic channel power.
Then, we transform $N^{2}$ time-of-flight measurement links into the target location. The Jacobian matrices of the $N$ BS measurement errors, evaluated at the true target position $\bf{t}$, can be expressed as $\mathbf{J}=\begin{bmatrix}\left(\frac{({\bf{b}}_{1}-\mathbf{t})^{T}}{\|{\bf{b}}_{1}-\mathbf{t}\|}+\frac{({\bf{b}}_{1}-\mathbf{t})^{T}}{\|{\bf{b}}_{1}-\mathbf{t}\|}\right)^{T}\;\cdots\;\left(\frac{({\bf{b}}_{N}-\mathbf{t})^{T}}{\|{\bf{b}}_{N}-\mathbf{t}\|}+\frac{({\bf{b}}_{N}-\mathbf{t})^{T}}{\|{\bf{b}}_{N}-\mathbf{t}\|}\right)^{T}\end{bmatrix}^{T}\in\mathbb{R}^{N^{2}\times 3}.$ Then, the Fisher information matrix (FIM) of sensing performance related to the target point is given by
$\mathbf{F}\left({\bf{t}}\right)=\mathbf{J}^{T}\mathbf{R}^{-1}\mathbf{J},\vspace{-1.5mm}$ (6)
where $\mathbf{R}^{-1}\!=\!\mathrm{diag}\left(\frac{1}{\|{\bf{b}}_{1}-\mathbf{t}\|^{2\beta}},\;\frac{1}{\|{\bf{b}}_{1}-\mathbf{t}\|^{\beta}\|{\bf{b}}_{2}-\mathbf{t}\|^{\beta}},\;\ldots,\;\frac{1}{\|{\bf{b}}_{N}-\mathbf{t}\|^{2\beta}}\right)$ $\in\mathbb{R}^{N^{2}\times N^{2}}.$ Since the transmit beamforming gain $G_{t}$ in the direction of the target is influenced by the randomness of the communication user channels in the joint waveform design (c.f. (2)), we perform an expected CRLB analysis based on the channel statistics. Given the locations of the ISAC BSs, the expected CRLB at a specific target point $\mathbf{t}$ is defined as
$\text{CRLB}({\bf{t}})={\rm{E}}_{G_{t}}\left[\text{trace}\left(\mathbf{F}\left({\bf{t}}\right)^{-1}\right)\right].\vspace{-1.5mm}$ (7)
@@
[…]
where $|\mathcal{A}|$ denotes the size (or volume) of the sensing area $\mathcal{A}$ in a $d$-dimensional space.
@@ II-C Cooperative Communication Model
We assume that the transmitters use non-coherent joint transmission, where the useful signals are combined by accumulating the transmit signal powers of cooperative BSs. Here, ${\bf{u}}=[x^{u},y^{u},z^{u}]^{T}\in\mathbb{R}^{3}$ denotes the user location. The signal received at the typical user is then given by
$y_{c}=$ $\underbrace{\sum\nolimits_{n=1}^{N}\|{\bf{b}}_{n}-{\bf{u}}\|^{-\frac{\alpha}{2}}\mathbf{h}_{n}^{H}\mathbf{W}_{n}\mathbf{s}_{1}}_{\text{collaborative intended signal}}$
$+\underbrace{\sum\nolimits_{{n\in\Phi_{I}}}\|{\bf{b}}_{n}-{\bf{u}}\|^{-\frac{\alpha}{2}}\mathbf{h}_{n}^{H}\mathbf{W}_{n}\mathbf{s}_{n}}_{\text{inter-cluster interference}}+{n_{c}},\vspace{-1.5mm}$ (9)
@@
${\rm{SNR}}_{c}=\frac{{\sum\limits_{n\in\Phi_{I}}{{g_{n}}{{\|{\bf{b}}_{n}-{\bf{u}}\|}^{-\alpha}}}}}{\sigma_{c}^{2}},\vspace{-1.5mm}$ (10)
where $g_{n}=p^{c}\left|\mathbf{h}_{n}^{H}\mathbf{w}_{n}^{c}\right|^{2}$ denotes the effective desired signals’ channel gain. The average data rate of users is given by
$R_{c}({\bf{u}})=\mathrm{E}_{g_{n}}[\log(1+\mathrm{SNR}_{c})].\vspace{-1.5mm}$ (11)
We then define the area communication rate over the communication user area $\mathcal{B}$ by $\frac{1}{{\cal{B}}}\int_{\bf{u}\in{|\cal{B}|}}{{R_{c}}\left(\bf{u}\right){\rm{d}}\bf{u}}$. Similarly, $|\mathcal{B}|$ denotes the size (or volume) of the communication user area ${\mathcal{B}}$.
@@ II-D Problem Formulation
where ${\bf{b}}_{n}$ denotes the position of the $n$th BS.
@@
In contrast to communication sum rate maximization, where the sum rate can theoretically grow unbounded as the SNR increases, fairness is difficult to ensure. The system tends to allocate more resources to users having better channel conditions, potentially leaving users with poor channels underserved or even entirely ignored. However, since the CRLB is strictly positive, minimizing the area CRLB can be heavily influenced by sub-regions having high CRLB values, even if the CRLB is near zero across most of the sensing area. Consequently, minimizing the area CRLB is inherently sensitive to worst-case locations. Explicitly, if a sub-region’s CRLB tends to infinity, so does the objective, forcing BS deployment to eliminate worse sensing points and thus yielding a naturally balanced performance over the area of interest without an explicit upper-bound constraint of $\max_{\mathbf{t}_{k}}\mathrm{CRLB}(\mathbf{t}_{k})\leq\mathrm{CRLB}^{\mathrm{th}}$.
@@ III-A Invariance Characteristics of Optimal Deployment
[…]
@@
${\bf{b}}_{n}^{\prime}-\mathbf{t}^{\prime}=({\bf{b}}_{n}+\mathbf{\Delta})-(\mathbf{t}+\mathbf{\Delta})={\bf{b}}_{n}-\mathbf{t}.\vspace{-1.5mm}$ (16)
Thus, relative positions remain unchanged, and the performance metric is displacement-invariant, i.e., $\mathrm{CRLB}(\mathbf{t}+\boldsymbol{\Delta})\;=\;\mathrm{CRLB}(\mathbf{t})$. If ${\bf{b}}_{n}^{*}$ is optimal for area $\mathcal{A}$, then ${\bf{b}}_{n}^{*}+\mathbf{\Delta}$ also achieves the same optimal metric $E^{*}$. By the method of contradiction, suppose $\{{\bf{b}}_{n}^{*}+\boldsymbol{\Delta}\}_{n=1}^{N}$ were not optimal on $\mathcal{A}_{\boldsymbol{\Delta}}$, i.e., a certain deployment achieves $E^{\prime}<E^{*}$ there; then displacing it back by $-\boldsymbol{\Delta}$ yields $E^{\prime}<E^{*}$ on $\mathcal{A}$, contradicting the optimality of $\{{\bf{b}}_{n}^{*}\}_{n=1}^{N}$. ∎
@@
[…]
@@
[…]
@@
[…]
@@
$\frac{\bigl{(}\mathbf{O}_{H}\,\mathbf{b}_{n}^{*}-\mathbf{O}_{H}\,\mathbf{t}\bigr{)}^{T}}{\bigl{\|}\mathbf{O}_{H}\,\mathbf{b}_{n}^{*}-\mathbf{O}_{H}\,\mathbf{t}\bigr{\|}}=\frac{(\mathbf{b}_{n}^{*}-\mathbf{t})^{T}\,\mathbf{O}_{H}^{T}}{\|\mathbf{b}_{n}^{*}-\mathbf{t}\|},\vspace{-1.5mm}$ (17)
@@
$\mathbf{F}(\mathbf{t}^{\prime})=(\mathbf{O}_{H}\,\mathbf{J}^{T})\,\mathbf{R}^{-1}\,(\mathbf{J}\,\mathbf{O}_{H}^{T})=\mathbf{O}_{H}\,\mathbf{F}(\mathbf{t})\,\mathbf{O}_{H}^{T},\vspace{-1.5mm}$ (18)
and therefore $(\mathbf{F}(\mathbf{t}^{\prime}))^{-1}=\mathbf{O}_{H}\,\mathbf{F}(\mathbf{t})^{-1}\,\mathbf{O}_{H}^{T}$. By the cyclic property of the trace, we have $\mathrm{trace}\bigl{(}(\mathbf{F}(\mathbf{t}^{\prime}))^{-1}\bigr{)}=\mathrm{trace}\bigl{(}\mathbf{F}(\mathbf{t})^{-1}\bigr{)}$, so the CRLB is invariant. Similar to the contradiction-based proof in Proposition 1, it can be shown that the reflected deployment attains the same minimum $\overline{\text{CRLB}}=E^{*}$ and it is optimal. ∎
The foregoing analysis has identified a set of invariance properties characterizing optimal BS deployments over a fixed-area region, which can facilitate low-complexity algorithm design, as detailed in Section III-C. In the next section, we extend this framework to variable-area scenarios and demonstrate how those invariances yield simple scaling laws that, in turn, lead to low-complexity algorithms for designing deployments that achieve optimal sensing coverage.

### v2

@@ II-B Cooperative Sensing Model
We aim to explore the optimal BS deployment method by examining time-of-flight based ranging measurements. The location of a certain target point is denoted as ${\bf{t}}\in\mathbb{R}^{3}$, where ${\bf{t}}\in\mathcal{A}$, and $\mathcal{A}$ represents the entire area of interest. The base-band equivalent of the signal reflected from target point $\bf{t}$} at receiver $m$ is represented as
@@
where $\beta\geq 2$ is the pathloss exponent between the serving BS and the target. [^3] For sensing performance analysis, we consider an open area of interest for target sensing where no blockage exists. This is because the radar sensing normally focuses on the light-of-sight (LoS) targets parameters and Non-LoS (NLoS) reflections may be neglected due to additional high pathloss of double-bouncing nature of radar. Furthermore, $\sigma$ denotes the radar cross section (RCS), $\tau_{n,m}$ is the propagation delay of the bistatic link spanning from BS $n$ to the target and then to BS $m$, while $\tilde{\tau}_{n,m}$ denotes the propagation delay of the direct link impinging from BS $n$ to BS $m$. Recognizing that a link-invariant RCS is an idealization in multistatic settings, we adopt an average effective $\sigma$ for tractable deployment optimization, consistent with [35]. In (3), ${\bf{H}}_{n,m}$ denotes the channel from BS $n$ to BS $m$. Finally, the term ${\bf{n}}(\tau)$ is the additive complex Gaussian noise having zero mean and covariance matrix $\sigma_{s}^{2}{\bf{I}}_{M_{\rm{r}}}$.
@@
$\mathbb{E}\left[\|\hat{\mathbf{t}}-\mathbf{t}\|^{2}\right]\geq\mathrm{Tr}\left(\mathbf{F}(\mathbf{t})^{-1}\right)\triangleq\mathrm{CRLB}(\mathbf{t}),\vskip 0.0pt$ (4)
where $\hat{{\bf{t}}}=\left[\hat{x}^{t},\hat{y}^{t},\hat{z}^{t}\right]^{T}$ represents the estimated location of the target. By applying matched filtering and reciprocal filtering, one can estimate the ToF of the signal between targets and BSs, and hence determine the propagation distance. Specifically, from transmitter $n$ to the target and then to receiver $m$, the term $\hat{d}_{nm}$ denotes the bistatic range corresponding to the path BS-$n$ $\to$ target $\to$ BS-$m$, which is given by
$\hat{d}_{nm}=d_{nm}+n^{t}_{nm},\vskip 0.0pt$ (5)
where $d_{nm}=\|\mathbf{b}_{n}-\mathbf{t}\|+\|\mathbf{b}_{m}-\mathbf{t}\|$ denotes the true distance, the measurement noise is $n^{t}_{nm}\sim\mathcal{N}\left(0,\eta_{nm}^{2}\right)$, $\eta_{nm}^{2}=\frac{3c^{2}\sigma_{s}^{2}}{8\pi^{2}G_{t}M_{r}B^{2}\gamma_{nm}}$, and $\gamma_{nm}=\bar{\sigma}\|\mathbf{b}_{n}-\mathbf{t}\|^{-\beta}\,\|\mathbf{b}_{m}-\mathbf{t}\|^{-\beta}$, the constant $\bar{\sigma}>0$ absorbs distance-invariant factors such as the average RCS. Here, $c$ denotes the speed of light, $B^{2}$ represents the effective squared bandwidth, $G_{t}$ is the transmit beamforming gain in the direction of the target, and $\gamma_{nm}$ represents the bistatic channel power.
Then, we transform $N^{2}$ time-of-flight measurement links into the target location. The Jacobian of the $N^{2}$ bistatic range measurements, evaluated at the true target position $\bf{t}$, [^4] In our system, while we may not have the exact target location, the estimated or predicted location of the target gleaned from previous sensing results can be utilized for performance analysis, such as in target tracking. can be expressed as
$\mathbf{J}(\mathbf{t})=\begin{bmatrix}\left(\frac{(\mathbf{t}-\mathbf{b}_{1})^{T}}{\|\mathbf{t}-\mathbf{b}_{1}\|}+\frac{(\mathbf{t}-\mathbf{b}_{1})^{T}}{\|\mathbf{t}-\mathbf{b}_{1}\|}\right)\\ \left(\frac{(\mathbf{t}-\mathbf{b}_{1})^{T}}{\|\mathbf{t}-\mathbf{b}_{1}\|}+\frac{(\mathbf{t}-\mathbf{b}_{2})^{T}}{\|\mathbf{t}-\mathbf{b}_{2}\|}\right)\\ \vdots\\ \left(\frac{(\mathbf{t}-\mathbf{b}_{n})^{T}}{\|\mathbf{t}-\mathbf{b}_{n}\|}+\frac{(\mathbf{t}-\mathbf{b}_{m})^{T}}{\|\mathbf{t}-\mathbf{b}_{m}\|}\right)\\ \vdots\\ \left(\frac{(\mathbf{t}-\mathbf{b}_{N})^{T}}{\|\mathbf{t}-\mathbf{b}_{N}\|}+\frac{(\mathbf{t}-\mathbf{b}_{N})^{T}}{\|\mathbf{t}-\mathbf{b}_{N}\|}\right)\end{bmatrix}\in\mathbb{R}^{N^{2}\times 3}.$
Then, the Fisher information matrix (FIM) of sensing performance related to the target point is given by
$\mathbf{F}\left({\bf{t}}\right)=\mathbf{J}^{T}\mathbf{R}^{-1}\mathbf{J},\vskip 0.0pt$ (6)
where $\mathbf{R}^{-1}=\mathrm{diag}\left(\eta_{nm}^{-2}\right)_{(n,m)}.$ Since the transmit beamforming gain $G_{t}$ in the direction of the target is influenced by the randomness of the communication user channels in the joint waveform design (c.f. (2)), we perform an expected CRLB analysis based on the channel statistics. Given the locations of the ISAC BSs, the expected CRLB at a specific target point $\mathbf{t}$ can be given by
$\text{CRLB}({\bf{t}})={\rm{E}}_{G_{t}}\left[\text{Tr}\left(\mathbf{F}\left({\bf{t}}\right)^{-1}\right)\right].\vskip 0.0pt$ (7)
@@
$\overline{\text{CRLB}}=\frac{1}{|\mathcal{A}|}\int_{\bf{t}\in{\mathcal{A}}}{{\rm{CRLB}}\left(\bf{t}\right)d{\bf{t}}},\vskip 0.0pt$ (8)
where $|\mathcal{A}|$ denotes the size (or volume) of the sensing area $\mathcal{A}$ in a $d$-dimensional space. As summarized in Table I, the individual-target CRLB is the pointwise CRLB on localization error for a single target. By contrast, the $d$-dimensional area CRLB averages $\mathrm{CRLB}(\mathbf{t})$ over a region $A\subset\mathbb{R}^{d}$ to assess localization performance across the entire area.
@@ II-C Cooperative Communication Model
We assume that the transmitters use non-coherent joint transmission, where the useful signals are combined by accumulating the transmit signal powers of cooperative BSs, i.e., all cooperating BSs transmit the same communication signal, with $\mathbf{s}_{n}=\left[s^{s}_{n},s^{c}\right]^{T}$. Here, ${\bf{u}}=[x^{u},y^{u},z^{u}]^{T}\in\mathbb{R}^{3}$ denotes the user location. The signal received at the typical user is then given by
$y_{c}=$ $\underbrace{\sum\nolimits_{n=1}^{N}\|{\bf{b}}_{n}-{\bf{u}}\|^{-\frac{\alpha}{2}}\mathbf{h}_{n}^{H}\mathbf{W}_{n}\mathbf{s}_{n}}_{\text{collaborative intended signal}}$
$+\underbrace{\sum\nolimits_{{n\in\Phi_{I}}}\|{\bf{b}}_{n}-{\bf{u}}\|^{-\frac{\alpha}{2}}\mathbf{h}_{n}^{H}\mathbf{W}_{n}\mathbf{s}_{n}}_{\text{inter-cluster interference}}+{n_{c}},\vskip 0.0pt$ (9)
@@
${\rm{SNR}}_{c}=\frac{{\sum\nolimits_{n=1}^{N}{{g_{n}}{{\|{\bf{b}}_{n}-{\bf{u}}\|}^{-\alpha}}}}}{\tilde{\sigma}_{c}^{2}},\vskip 0.0pt$ (10)
where $\left|\mathbf{h}_{n}^{H}\mathbf{w}_{n}^{c}\right|^{2}$ denotes the effective desired signals’ channel gain, and $\tilde{\sigma}_{c}^{2}$ denotes the effective noise variance, including interference. The average data rate of users is given by
$R_{c}({\bf{u}})=\mathrm{E}_{g_{n}}[\log_{2}(1+\mathrm{SNR}_{c})].\vskip 0.0pt$ (11)
We then define the area communication rate over the communication user area $\mathcal{B}$ by $\frac{1}{{|\mathcal{B}|}}\int_{\bf{u}\in{\mathcal{B}}}{{R_{c}}\left(\bf{u}\right){\rm{d}}\bf{u}}$. Similarly, $|\mathcal{B}|$ denotes the size (or volume) of the communication user area $\mathcal{B}$}.
@@ II-D Problem Formulation
where ${\bf{b}}_{n}$ denotes the position of the $n$th BS. [^5] While our study treats static BS placement, the framework naturally extends to mobile BSs (e.g., drones) by re-optimizing time-varying locations on a receding horizon using online user and target trajectory estimates, subject to mobility and energy constraints, to minimize a time-averaged area-CRLB.
@@
In contrast to communication sum rate maximization, where the sum rate can theoretically grow unbounded as the SNR increases, fairness is difficult to ensure. The system tends to allocate more resources to users having better channel conditions, potentially leaving users with poor channels underserved or even entirely ignored. However, since the CRLB is strictly positive, minimizing the area CRLB can be heavily influenced by sub-regions having high CRLB values, even if the CRLB is near zero across most of the sensing area. This characteristic makes the optimization highly sensitive to regions with poor performance, compelling BS deployment to prioritize these regions for improvement. Consequently, minimizing the area CRLB is inherently sensitive to worst-case locations. Explicitly, if a sub-region’s CRLB tends to infinity, so does the objective, forcing BS deployment to eliminate worse sensing points and thus yielding a naturally balanced performance over the area of interest without an explicit upper-bound constraint of $\max_{\mathbf{t}_{k}}\mathrm{CRLB}(\mathbf{t}_{k})\leq\mathrm{CRLB}^{\mathrm{th}}$.
@@ III-A Invariance Characteristics of Optimal Deployment
Invariance to displacement. Let $\mathcal{A}\subset\mathbb{R}^{d}$ be the sensing area in problem $(\mathrm{P1.1})$, and let $\{{\bf{b}}_{n}^{*}\}_{n=1}^{N}$ be an optimal BS deployment attaining the minimum $\overline{\text{CRLB}}$ $E^{*}$. For any displacement vector $\boldsymbol{\Delta}\in\mathbb{R}^{d}$, define the shifted area $\mathcal{A}_{\boldsymbol{\Delta}}=\mathcal{A}+\boldsymbol{\Delta}$. Then, $\{{\bf{b}}_{n}^{*}+\boldsymbol{\Delta}\}_{n=1}^{N}$ is an optimal solution of $(\mathrm{P1.1})$ posed over $\mathcal{A}_{\boldsymbol{\Delta}}$, achieving the same objective value $E^{*}$.
@@
${\bf{b}}_{n}^{\prime}-\mathbf{t}^{\prime}=({\bf{b}}_{n}+\mathbf{\Delta})-(\mathbf{t}+\mathbf{\Delta})={\bf{b}}_{n}-\mathbf{t}.\vskip 0.0pt$ (16)
Thus, relative positions remain unchanged, and the performance metric is displacement-invariant, i.e., $\mathrm{CRLB}(\mathbf{t}+\boldsymbol{\Delta})\;=\;\mathrm{CRLB}(\mathbf{t})$. If ${\bf{b}}_{n}^{*}$ is optimal for area $\mathcal{A}$, then ${\bf{b}}_{n}^{*}+\mathbf{\Delta}$ also achieves the same optimal metric $E^{*}$. By the method of contradiction, suppose $\{{\bf{b}}_{n}^{*}+\boldsymbol{\Delta}\}_{n=1}^{N}$ were not optimal on $\mathcal{A}_{\boldsymbol{\Delta}}$, i.e., a certain deployment achieves $E^{\prime}<E^{*}$ there; then displacing it back by $\boldsymbol{\Delta}$ yields $E^{\prime}<E^{*}$ on $\mathcal{A}$, contradicting the optimality of $\{{\bf{b}}_{n}^{*}\}_{n=1}^{N}$. ∎
@@
Invariance to Rotation. Let $\mathcal{A}\subset\mathbb{R}^{d}$ be the sensing area in problem $(\mathrm{P1.1})$. Suppose $\{{\bf{b}}_{n}^{*}\}_{n=1}^{N}$ is an optimal BS deployment for $(\mathrm{P1.1})$, attaining the minimum $\overline{\text{CRLB}}$ $E^{*}$. Let ${\bf{R}}_{\varphi}\in\mathbb{R}^{d\times d}$ be any arbitrary rotation matrix, i.e. ${\bf{R}}_{\varphi}{\bf{R}}_{\varphi}^{T}=\mathbf{I},\det({\bf{R}}_{\varphi})=1,$ and define the rotated regions $\mathcal{A}_{{\bf{R}}_{\varphi}}={\bf{R}}_{\varphi}\mathcal{A}$. Then $\{{\bf{R}}_{\varphi}\,{\bf{b}}_{n}^{*}\}_{n=1}^{N}$ is an optimal solution of $(\mathrm{P1.1})$ over $\mathcal{A}_{{\bf{R}}_{\varphi}}$, achieving the same objective value $E^{*}$.
@@
For any ${\bf{b}},\mathbf{t}\in\mathbb{R}^{d}$, $\|{\bf{R}}_{\varphi}{\bf{b}}-{\bf{R}}_{\varphi}\mathbf{t}\|=\|{\bf{b}}-\mathbf{t}\|,$ we have $\mathrm{CRLB}({\bf{R}}_{\varphi}\mathbf{t})=\mathrm{CRLB}(\mathbf{t}).$ Therefore, it follows that $\frac{1}{|\mathcal{A}_{{\bf{R}}_{\varphi}}|}\int_{\mathcal{A}_{{\bf{R}}_{\varphi}}}\mathrm{CRLB}(\mathbf{x})\mathrm{d}\mathbf{x}=\frac{1}{|\mathcal{A}|}\int_{\mathcal{A}}\mathrm{CRLB}(\mathbf{t})\mathrm{d}\mathbf{t}=E^{*}.$ If a deployment on $\mathcal{A}_{{\bf{R}}_{\varphi}}$ achieved an objective less than $E^{*}$, rotating it back by ${\bf{R}}_{\varphi}^{T}$ would yield a deployment on $\mathcal{A}$ having an objective below $E^{*}$, contradicting the optimality of $\{{\bf{b}}_{n}^{*}\}$. Hence $\{{\bf{b}}_{n}^{\prime}\}$ is optimal on $\mathcal{A}_{{\bf{R}}_{\varphi}}$. Similar to the proof in Proposition 1, it can be shown that $\{{\bf{R}}_{\varphi}\,{\bf{b}}_{n}^{*}\}_{n=1}^{N}$ is an optimal solution of $(\mathrm{P1.1})$ over $\mathcal{A}_{{\bf{R}}_{\varphi}}$. ∎
@@
Invariance under Symmetric Projection. Let $\mathcal{A}\subset\mathbb{R}^{d}$ be the sensing region in problem $(\mathrm{P1.1})$. Let $\mathbf{O}_{H}\in\mathbb{R}^{d\times d},\;\mathbf{O}_{H}\mathbf{O}_{H}^{T}=\mathbf{I},\;\det(\mathbf{O}_{H})=-1,$ be the symmetric projection matrix across some hyperplane $H$, and define its symmetric projection region $\mathcal{A}^{\prime}\;=\;\mathbf{O}_{H}\,\mathcal{A}.$ If $\{\mathbf{b}_{n}^{*}\}_{n=1}^{N}$ is an optimal BS deployment on $\mathcal{A}$ achieving $\overline{\mathrm{CRLB}}=E^{*}$, then the reflected deployment $\{\mathbf{O}_{H}\,\mathbf{b}_{n}^{*}\}_{n=1}^{N}$ is optimal on $\mathcal{A}^{\prime}$, achieving the same $\overline{\mathrm{CRLB}}=E^{*}$.
@@
$\frac{\bigl(\mathbf{O}_{H}\,\mathbf{b}_{n}^{*}-\mathbf{O}_{H}\,\mathbf{t}\bigr)^{T}}{\bigl\|\mathbf{O}_{H}\,\mathbf{b}_{n}^{*}-\mathbf{O}_{H}\,\mathbf{t}\bigr\|}=\frac{(\mathbf{b}_{n}^{*}-\mathbf{t})^{T}\,\mathbf{O}_{H}^{T}}{\|\mathbf{b}_{n}^{*}-\mathbf{t}\|},\vskip 0.0pt$ (17)
@@
$\mathbf{F}(\mathbf{t}^{\prime})=(\mathbf{O}_{H}\,\mathbf{J}^{T})\,\mathbf{R}^{-1}\,(\mathbf{J}\,\mathbf{O}_{H}^{T})=\mathbf{O}_{H}\,\mathbf{F}(\mathbf{t})\,\mathbf{O}_{H}^{T},\vskip 0.0pt$ (18)
and therefore $(\mathbf{F}(\mathbf{t}^{\prime}))^{-1}=\mathbf{O}_{H}\,\mathbf{F}(\mathbf{t})^{-1}\,\mathbf{O}_{H}^{T}$. By the cyclic property of the trace, we have $\mathrm{Tr}\bigl((\mathbf{F}(\mathbf{t}^{\prime}))^{-1}\bigr)=\mathrm{Tr}\bigl(\mathbf{F}(\mathbf{t})^{-1}\bigr)$, so the CRLB is invariant. Similar to the contradiction-based proof in Proposition 1, it can be shown that the reflected deployment attains the same minimum $\overline{\text{CRLB}}=E^{*}$ and it is optimal. ∎
The foregoing analysis has identified a set of invariance properties characterizing optimal BS deployments over a fixed‐area region, which can facilitate low-complexity algorithm design, as detailed in Section III-C. In the next section, we extend this framework to variable‐area scenarios and demonstrate how those invariances yield simple scaling laws that, in turn, lead to low‐complexity algorithms for designing deployments that achieve optimal sensing coverage.